OpenStudy (anonymous):

Given the trinomial 2x^2+4x-2, what is the value of the discriminant? A.0 B.10 C.32 D.36

OpenStudy (anonymous):

\[ 2x^2+4x-2 \] comparing it with \[ ax^2+bx+c\] we find a= 2, b= 4 & c=2 now discriminant is given as: \[b^2-4ac= 4^2-4 \times \times 2\times 2 = 16-16=0\] Hence option A.0 is the correct answer. @AshleighW

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

Sure how?...sorry I just made a account like 10 mins ago lol

OpenStudy (anonymous):

@dpasinghCan you help me with 4 more questions...ill give you medals for all?

OpenStudy (anonymous):

@AshleighW sure, go aead

OpenStudy (anonymous):

Given the trinomial 2x^2+4x+4,predict the type of solutions? A.2 rational solutions B.1 rational solution C.2 irrational solutions D.2 complex solutions

OpenStudy (anonymous):

Select one of the factors x^3y^2+8xy^2-5x^2-40 A.(xy^2+5) B.(x^2+4) C.(xy^2-5) D.(x^2-8)

OpenStudy (anonymous):

\[2x^2+4x+4 \rightarrow ax^2+bx+c \rightarrow a= 2, b= 4, c= 4\] now let us find its discriminant i.e. \[b^2-4ac = 4^2-4 \times 2 \times 4= 16- 32=-16 < 0 \] Thus It will have 2 complex solutions. Hence option D.2 complex solutions is the correct Answer. @@AshleighW

OpenStudy (anonymous):

@AshleighW

OpenStudy (anonymous):

Select one of the factors of 3x^2+10x+3 A.(3x+1) B.(3x-1) C.3x+3) D.None of the above

OpenStudy (anonymous):

\[x^3y^2+8xy^2-5x^2-40 = xy^2( x^2+8)-5(x^2+8) =(x^2+8)(8xy^2-5)\] Hence \[(x^2+8)(xy^2-5) \] are the required factors. Hence option C.(xy^2-5) is the correct answer. @AshleighW

OpenStudy (anonymous):

And finally.. Select one of the factors of 3x^2+4x-4 A.(x+4) B.(x-2) C.(3x+2) D.(3x-2)

OpenStudy (anonymous):

Thanks if you could please answer those last two:)